# Preserve or Modify? Context-Aware Evaluation for Balancing Preservation and Modification in Text-Guided Image Editing

### v2

@@ Abstract
The development of vision-language and generative models has significantly advanced text-guided image editing, which seeks the preservation of core elements in the source image while implementing modifications based on the target text. However, existing metrics have a context-blindness problem, indiscriminately applying the same evaluation criteria on completely different pairs of source image and target text, biasing towards either modification or preservation. Directional CLIP similarity, the only metric that considers both source image and target text, is also biased towards modification aspects and attends to irrelevant editing regions of the image. We propose AugCLIP, a context-aware metric that adaptively coordinates preservation and modification aspects, depending on the specific context of a given source image and target text. This is done by deriving the CLIP representation of an ideally edited image, that preserves the source image with necessary modifications to align with target text. More specifically, using a multi-modal large language model, AugCLIP augments the textual descriptions of the source and target, then calculates a modification vector through a hyperplane that separates source and target attributes in CLIP space. Extensive experiments on five benchmark datasets, encompassing a diverse range of editing scenarios, show that AugCLIP aligns remarkably well with human evaluation standards, outperforming existing metrics. The code will be open-sourced for community use.
\$*\$\$*\$footnotetext: Equal contribution\$$\dagger$\$$\dagger$\$footnotetext: Corresponding author
@@ 1 Introduction
Building on advancements in vision-language models (Radford et al., 2021; Li et al., 2022; Geng et al., 2023), recent generative models (Kawar et al., 2022; Brooks et al., 2022; Hertz et al., 2022) have been widely utilized as creative tools for image editing via text instructions. Text-guided image editing models enable the modification of images in response to textual guidance, ensuring that changes are aligned with the provided instructions. The primary objective of these models is to apply the necessary modifications guided by the target text while preserving most of the source image.
@@
Despite the remarkable advancements in editing models, there has been a lack of rigorous evaluation methods, tailored specifically for text-guided image editing. Consequently, most studies (Hertz et al., 2023; Basu et al., 2023; Gal et al., 2022; Kim & Ye, 2021; Brooks et al., 2022; Ruiz et al., 2023; Kocasari et al., 2022) have heavily relied on human evaluation, which provides balanced consideration of preservation and modification aspects. However, as it is costly and impractical for real-world applications, researchers have adapted automatic evaluation metrics (Heusel et al., 2017; Zhang et al., 2018; Caron et al., 2021; Hessel et al., 2021) originally designed for other vision tasks, such as image generation or captioning.
@@
Among existing metrics, directional CLIP similarity (Gal et al., 2022) is unique because it provides a method to consider both the source image and target text in the evaluation. Even this metric suffers from a context-blindness problem, using a fixed standard without considering the specific relationship between the source image and target text. We show that directional CLIP similarity tends to prioritize modification over preservation, and often focuses on peripheral parts rather than regions that are edited as guided by target text, in Sec. 3. These observations underscore the need for a context-aware metric to dynamically adjust the evaluation standard, balancing both aspects in response to diverse editing contexts.
@@ 2 Background on Existing Metrics
FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), and L2 distance measure feature-level distance between images. These metrics are primarily used in image generation tasks to compare the distributions of real and generated images. When applied to editing scenarios, they assess the distance between the source and edited images, completely ignoring the target text guiding the editing process. Segmentation consistency (SC) (Kim & Ye, 2021) measures the structural change. Additionally, DINO similarity (DINO) and CLIP-I evaluate the semantic similarity between the source and edited images within the DINO (Caron et al., 2021) and CLIP (Radford et al., 2021) embedding spaces, also focusing solely on preservation.
@@
To assess how well an edited image aligns with the target text, the primary approach involves using a pre-trained multi-modal model (Radford et al., 2021). CLIPScore (CLIP-T) (Hessel et al., 2021) is the most widely adopted metric in this category, measuring the similarity between the edited image and the target text. This metric completely ignores the source image that should be considered for the preservation aspect.
@@
Directional CLIP similarity ($\text{CLIP}_{\text{dir}}$) (Gal et al., 2022) is unique among existing metrics as it assesses both preservation and modification aspects by evaluating directional alignment between the original and edited images concerning the source and target text.
@@
Metrics Used in Text-to-Image Generation
While text-to-image generation tasks benefit from a variety of evaluation metrics utilizing pre-trained multi-modal large language models (Hu et al., 2023; Li et al., 2024b; Lu et al., 2024; Huang et al., 2023; Cho et al., 2024) or VQA models (Hu et al., 2023; Ghosh et al., 2024), these metrics focus solely on prompt faithfulness. In contrast, image editing scenarios require a balance of image preservation and prompt adherence, making these metrics irrelevant to the evaluation of editing models.
@@ 3.1 Combination of Preservation- and Modification-Centric Metrics
We explore whether combining preservation- and modification-centric metrics can improve alignment with human judgments across two benchmark datasets: EditVal (Basu et al., 2023) and CelebA (Liu et al., 2015) (Results on the other three benchmark datasets are deferred to the Appendix). To do so, we combine CLIP-T, a modification-centric metric, and preservation-centric metrics (LPIPS, DINO similarity, CLIP-I) with varying interpolation values $\gamma\in[0,1]$. Note that both metrics are scaled to the same range before interpolation, ensuring the interpolation value $\gamma$ is properly reflected in the final combined score.
@@ 3.2 Directional CLIP Similarity
Directional CLIP similarity overly emphasizes the alignment with the target text while neglecting the preservation of the source image. Despite attempting to incorporate preservation by subtracting $T_{\text{src}}$ from $T_{\text{trg}}$, $\mathrm{CLIP}_{\text{dir}}$ frequently favors excessively modified images over ground truth ideally edited images as shown in samples in Fig. 3(a). Specifically, $\mathrm{CLIP}_{\text{dir}}$ metric fails on $63\%$ of the 100 editing cases in TEdBench (Kawar et al., 2022) and $35.1\%$ of the 1053 cases in MagicBrush (Zhang et al., 2024) (Details in Sec. 5). This problem happens since $\mathrm{CLIP}_{\text{dir}}$ is designed to assume that a well-edited image should primarily adhere to the target text. Thus, it blindly favors the visual features of the target text, leading to unreliable results.
@@
$\mathrm{CLIP}_{\text{dir}}$ often focuses on peripheral or unmodified local areas of the edited image, overlooking the specific areas that are edited by the text. Using the relevancy map (Chefer et al., 2021), ${\bm{R}}(\Delta I;\Delta T)$ [^1] The original relevancy map ${\bm{R}}$ utilizes an image-text pair. To adapt it to $\mathrm{CLIP}_{\text{dir}}$, we subtract two relevancy maps. Details are in the Appendix., to visualize attention on image regions corresponding to the text direction, we observe that $\mathrm{CLIP}_{\text{dir}}$ frequently fails to concentrate on the image regions that are directly relevant to the target text as shown in Fig. 3(b). The figure also shows that adding explicit visual details guides $\mathrm{CLIP}_{\text{dir}}$ to attend to edited regions properly.
@@ 4.1 Extracting and Processing Attributes
As described in Fig. 4(a), we extract visual descriptions about attributes of the source image and target text using a state-of-the-art multi-modal large language model (MLLM), GPT-4V (OpenAI, 2023). From the source image, we extract visual features that can be the object of preservation. Likewise, from the target text, we extract visual features that can be the object of modification. The extraction process is instructed by five annotated in-context examples, which are provided in the Appendix.
@@ 4.2 Deriving the Ideal Editing in CLIP Space
Table 1: Benchmark Datasets and Editing Models. Benchmark dataset type that corresponds to different editing scenarios. For each dataset, we use different types of editing models that are tailored for each scenario.
$\dagger$ Kwon et al. (2022); Kim et al. (2022); Kim & Ye (2021); Patashnik et al. (2021)
$\ddagger$ Hertz et al. (2023); Couairon et al. (2022); Brooks et al. (2022); Hertz et al. (2022); Kawar et al. (2022)
$\ast$ Li et al. (2024a); Kumari et al. (2023); Wei et al. (2023)
@@
Editing Models $\dagger$ $\ddagger$ $\ddagger$ $\ddagger$ $\ast$
@@ 5 Experiments
For our experiments, we employ a pre-trained CLIP-ViT 16/B model for CLIP-based metrics. Source and target attributes are generated using GPT-4V (OpenAI, 2023). Further details on prompting the source and target attributes are in the Appendix.
@@
We evaluate AugCLIP and existing metrics across several text-guided image editing benchmarks, including TEdBench (Kawar et al., 2022), EditVal (Basu et al., 2023), MagicBrush (Zhang et al., 2024), DreamBooth (Ruiz et al., 2023), and CelebA Liu et al. (2015). Each benchmark dataset represents varying editing scenarios as in Tab. 1.
For each benchmark dataset, we employ multiple editing models (Kwon et al., 2022; Kim et al., 2022; Kim & Ye, 2021; Patashnik et al., 2021; Hertz et al., 2023; Couairon et al., 2022; Brooks et al., 2022; Hertz et al., 2022; Li et al., 2024a; Kumari et al., 2023; Wei et al., 2023) as specified in Tab. 1. For CelebA, which require facial attribute editing, we employ domain-specific editing models pre-trained on facial dataset. For EditVal, TEdBench, and MagicBrush, we employ generic-purpose editing models. For DreamBooth, we use presonalized generation models. More details on benchmark datasets are provided in the Appendix.
@@ 5.1 Correlation with Human Judgment
As shown in Tab. 3, AugCLIP, the context-aware metric that balances preservation and modification, is superior over all previous metrics across diverse benchmark datasets. A high level of alignment between preservation-centric (L2, LPIPS, DINO, SC, CLIP-I) and human judgment is expected because these benchmarks do not significantly deviate from the source image.
@@ 6 Analysis
Pos. Add Obj. repl. Alter Parts Background Texture Color Action Style
@@
AugCLIP 1.0 0.75 0.838 1.0 0.742 1.0 1.0 0.647
@@
Description Randomness
Using GPT-4V for description extraction introduces slight randomness across different seeds. Tests with five seeds show low variance ($0.0197$) in human alignment score (${\bm{s}}_{\text{2AFC}}$), indicating that the description generation process does not affect the scoring. We also demonstrate in the Appendix that extracted attributes with different seeds do not vary.
@@
Our method, AugCLIP, adds additional computation time compared to $\mathrm{CLIP}_{\text{dir}}$. It requires extracting attributes via MLLM, and then fitting a hyperplane to distinguish source and target attributes. This process takes around 12.3 seconds for attribute generation and 0.15 seconds for score computation. We will provide extracted attributes for all benchmark datasets to reduce this overhead, facilitating the evaluation process for future usage.
@@ A.3 Optimization Objective
Additionally, we visualize that CLIP features of source and target attributes can be separated by a linear hyperplane in 2D projected space using UMAP (McInnes et al., 2018) in Fig. 8, in which randomly chosen subset of TEdBench samples are plotted. ‘S’ represents source attributes encoded into CLIP, while ‘T’ represents the target attributes. The line signifies the linear hyperplane $g(x)=\mathbf{W}^{T}x+b=0$ that separates the two classes, which are source and target. In AugCLIP, the linear hyperplane $g(x)=0$ is $d$-dimensional following the original dimension of CLIP, but to visualize in Fig. 8, the dimension shrinks into $d=2$ by UMAP fitting.
@@ C.1 Assets
Category Asset URL
Benchmarks CelebA (Liu et al., 2015) https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html
TedBench (Kawar et al., 2022) https://github.com/imagic-editing/imagic-editing.github.io/tree/main/tedbench
EditVal (Basu et al., 2023) https://github.com/deep-ml-research/editval_code
DreamBooth (Ruiz et al., 2023) https://github.com/google/dreambooth
MagicBrush (Zhang et al., 2024) https://github.com/OSU-NLP-Group/MagicBrush
Editing Models TEdBench EditVal MagicBrush InstructPix2Pix (Brooks et al., 2022) https://github.com/timothybrooks/instruct-pix2pix
DiffEdit (Couairon et al., 2022) https://github.com/Xiang-cd/DiffEdit-stable-diffusion.git
Prompt-to-Prompt (Hertz et al., 2022) https://github.com/google/prompt-to-prompt.git
DDS (Hertz et al., 2023) https://github.com/google/prompt-to-prompt/blob/main/DDS_zeroshot.ipynb
Imagic (Kawar et al., 2022) https://github.com/huggingface/diffusers/tree/main/examples/community#imagic-stable-diffusion
CelebA DiffusionCLIP (Kim & Ye, 2021) https://github.com/gwang-kim/DiffusionCLIP
Multi2One (Kim et al., 2022) https://github.com/akatigre/multi2one
Asyrp (Kwon et al., 2022) https://github.com/kwonminki/Asyrp_official
StyleCLIP (Patashnik et al., 2021) https://github.com/orpatashnik/StyleCLIP
DreamBooth Custom Diffusion  (Kumari et al., 2023) https://github.com/adobe-research/custom-diffusion
BLIP-Diffusion (Li et al., 2024a) https://github.com/salesforce/LAVIS/tree/main/projects/blip-diffusion
ELITE (Wei et al., 2023) https://github.com/csyxwei/ELITE
Metrics CLIP (Radford et al., 2021) https://github.com/openai/CLIP
LPIPS (Zhang et al., 2018) https://pypi.org/project/lpips/
SC (Kim & Ye, 2021) Implemented by the Authors

### v3

@@ Abstract
The development of vision-language and generative models has significantly advanced text-guided image editing, which seeks the preservation of core elements in the source image while implementing modifications based on the target text. However, existing metrics have a context-blindness problem, indiscriminately applying the same evaluation criteria on completely different pairs of source image and target text, biasing towards either modification or preservation. Directional CLIP similarity, the only metric that considers both source image and target text, is also biased towards modification aspects and attends to irrelevant editing regions of the image. We propose AugCLIP, a context-aware metric that adaptively coordinates preservation and modification aspects, depending on the specific context of a given source image and target text. This is done by deriving the CLIP representation of an ideally edited image, that preserves the source image with necessary modifications to align with target text. More specifically, using a multi-modal large language model, AugCLIP augments the textual descriptions of the source and target, then calculates a modification vector through a hyperplane that separates source and target attributes in CLIP space. Extensive experiments on five benchmark datasets, encompassing a diverse range of editing scenarios, show that AugCLIP aligns remarkably well with human evaluation standards, outperforming existing metrics. The code is available at https://github.com/augclip/augclip_eval.
@@ 1 Introduction
Building on advancements in vision-language models [24, 19, 7], recent generative models [12, 2, 8] have been widely utilized as creative tools for image editing via text instructions. Text-guided image editing models enable the modification of images in response to textual guidance, ensuring that changes are aligned with the provided instructions. The primary objective of these models is to apply the necessary modifications guided by the target text while preserving most of the source image.
@@
Despite the remarkable advancements in editing models, there has been a lack of rigorous evaluation methods, tailored specifically for text-guided image editing. Consequently, most studies [9, 1, 6, 13, 2, 26, 15] have heavily relied on human evaluation, which provides balanced consideration of preservation and modification aspects. However, as it is costly and impractical for real-world applications, researchers have adapted automatic evaluation metrics [11, 29, 3, 10] originally designed for other vision tasks, such as image generation or captioning.
@@
Among existing metrics, directional CLIP similarity [6] is unique because it provides a method to consider both the source image and target text in the evaluation. Even this metric suffers from a context-blindness problem, using a fixed standard without considering the specific relationship between the source image and target text. We show that directional CLIP similarity tends to prioritize modification over preservation, and often focuses on peripheral parts rather than regions that are edited as guided by target text, in Sec. 3. These observations underscore the need for a context-aware metric to dynamically adjust the evaluation standard, balancing both aspects in response to diverse editing contexts.
@@ 2 Background on Existing Metrics
FID [11], LPIPS [29], and L2 distance measure feature-level distance between images. These metrics are primarily used in image generation tasks to compare the distributions of real and generated images. When applied to editing scenarios, they assess the distance between the source and edited images, completely ignoring the target text guiding the editing process. Segmentation consistency (SC) [13] measures the structural change. Additionally, DINO similarity (DINO) and CLIP-I evaluate the semantic similarity between the source and edited images within the DINO [3] and CLIP [24] embedding spaces, also focusing solely on preservation.
@@
To assess how well an edited image aligns with the target text, the primary approach involves using a pre-trained multi-modal model [24]. CLIPScore (CLIP-T) [10] is the most widely adopted metric in this category, measuring the similarity between the edited image and the target text. This metric completely ignores the source image that should be considered for the preservation aspect.
@@
Directional CLIP similarity ($\text{CLIP}_{\text{dir}}$) [6] is unique among existing metrics as it assesses both preservation and modification aspects by evaluating directional alignment between the original and edited images concerning the source and target text.
@@ 3.1 Combination of Preservation- and Modification-Centric Metrics
We explore whether combining preservation- and modification-centric metrics can improve alignment with human judgments across two benchmark datasets: EditVal [1] and CelebA [20] (Results on the other three benchmark datasets are deferred to the Appendix). To do so, we combine CLIP-T, a modification-centric metric, and preservation-centric metrics (LPIPS, DINO similarity, CLIP-I) with varying interpolation values $\gamma\in[0,1]$. Note that both metrics are scaled to the same range before interpolation, ensuring the interpolation value $\gamma$ is properly reflected in the final combined score.
@@ 3.2 Directional CLIP Similarity
Directional CLIP similarity overly emphasizes the alignment with the target text while neglecting the preservation of the source image. Despite attempting to incorporate preservation by subtracting $T_{\text{src}}$ from $T_{\text{trg}}$, $\mathrm{CLIP}_{\text{dir}}$ frequently favors excessively modified images over ground truth ideally edited images as shown in samples in Fig. 3(a). Specifically, $\mathrm{CLIP}_{\text{dir}}$ metric fails on $63\%$ of the 100 editing cases in TEdBench [12] and $35.1\%$ of the 1053 cases in MagicBrush [28] (Details in Sec. 5). This problem happens since $\mathrm{CLIP}_{\text{dir}}$ is designed to assume that a well-edited image should primarily adhere to the target text. Thus, it blindly favors the visual features of the target text, leading to unreliable results.
@@
$\mathrm{CLIP}_{\text{dir}}$ often focuses on peripheral or unmodified local areas of the edited image, overlooking the specific areas that are edited by the text. Using the relevancy map [4], ${\bm{R}}(\Delta I;\Delta T)$ [^1] The original relevancy map ${\bm{R}}$ utilizes an image-text pair. To adapt it to $\mathrm{CLIP}_{\text{dir}}$, we subtract two relevancy maps. Details are in the Appendix., to visualize attention on image regions corresponding to the text direction, we observe that $\mathrm{CLIP}_{\text{dir}}$ frequently fails to concentrate on the image regions that are directly relevant to the target text as shown in Fig. 3(b). The figure also shows that adding explicit visual details guides $\mathrm{CLIP}_{\text{dir}}$ to attend to edited regions properly.
@@ 4.1 Extracting and Processing Attributes
As described in Fig. 4(a), we extract visual descriptions about attributes of the source image and target text using a state-of-the-art multi-modal large language model (MLLM), GPT-4V [22]. From the source image, we extract visual features that can be the object of preservation. Likewise, from the target text, we extract visual features that can be the object of modification. The extraction process is instructed by five annotated in-context examples, which are provided in the Appendix.
@@ 4.2 Deriving the Ideal Editing in CLIP Space
Table 1: Benchmark Datasets and Editing Models Used for Evaluating the Performance of Metrics.
@@
Editing Models [17, 14, 13, 23] [9, 5, 2, 8, 12] [9, 5, 2, 8, 12] [9, 5, 2, 8, 12] [18, 16, 27]
@@ 5 Experiments
For our experiments, we employ a pre-trained CLIP-ViT 16/B model for CLIP-based metrics. Source and target attributes are generated using GPT-4V [22]. Further details on prompting the source and target attributes are in the Appendix.
@@
We evaluate AugCLIP and existing metrics across several text-guided image editing benchmarks, including TEdBench [12], EditVal [1], MagicBrush [28], DreamBooth [26], and CelebA [20]. Each benchmark dataset represents varying editing scenarios as in Tab. 1. For each benchmark dataset, we employ multiple editing models [17, 14, 13, 23, 9, 5, 2, 8, 18, 16, 27] as specified in Tab. 1 for suitable type of benchmark dataset. Details on benchmark datasets are provided in the Appendix.
@@ 5.1 Correlation with Human Judgment
As shown in Tab. 3, AugCLIP, the context-aware metric that balances preservation and modification, is superior over all previous metrics across diverse benchmark datasets. A high level of alignment between preservation-centric (L2, LPIPS, DINO, SC, CLIP-I) and human judgment is expected because these benchmarks do not significantly deviate from the source image. Moreover, per-model evaluations are presented alongside human evaluations in Tab. 4. Unlike other metrics, AugCLIP perfectly aligns with human rankings on each dataset.
@@
Table 4: Rank of various text-guided editing models. Highlighted in green if a metric aligns with human evaluations; otherwise, highlighted in red.
Dataset Models Rank $\mathrm{CLIP}_{\text{dir}}$uparrow LPIPS $\downarrow$ AugCLIP $\uparrow$ Human $\uparrow$
DreamBooth ELITE 1 0.1132 2 71.38 2 0.7642 1 0.8478
BlipDiffusion 2 0.0836 3 70.88 1 0.7579 2 0.6525
CustomDiffusion 3 0.1348 1 73.84 3 0.6156 3 0.0263
EditVal P2P 1 0.1771 3 15.04 1 0.8521 1 0.6133
InstructPix2Pix 2 0.1774 2 25.75 3 0.8242 2 0.4855
DiffEdit 3 0.2272 1 20.41 2 0.8155 3 0.3214
CelebA StyleCLIP 1 0.0376 2 27.04 1 0.8484 1 0.6831
Multi2One 2 0.0414 1 27.95 2 0.8152 2 0.5469
Asyrp 3 -0.0001 3 36.98 3 0.7750 3 0.3197
@@ 6 Analysis
Pos. Add Obj. repl. Alter Parts Background
@@
AugCLIP 1.0 0.75 0.838 1.0
Texture Color Action Style
[…]
AugCLIP 0.742 1.0 1.0 0.647
Design of Weighting
To support the rationale behind weighting, we design an experiment [^2] Two attributes are compared in terms of relative importance in the same context. A total of 10 participants responded in 2AFC format on 40 editing scenarios from the DreamBooth dataset. that measures pearson correlation with human evaluation on attribute importance, reported in Tab. 7. Compared to the naïve weighting strategy mimicking $\mathrm{CLIP}_{\text{dir}}$ by measuring cosine similarity between the attribute and the text direction $\Delta T$ (Eq. 1), AugCLIP weighting (Eq. 6) demonstrates impressive alignment with human evaluations.
Table 7: Pearson correlation between weighting strategies and human evaluations on attribute importance.
Attribute Importance Source Target
Naïve Weighting w/ $\Delta T$ 0.394 0.637
AugCLIP Weighting (Eq. 6) 0.794 0.883
@@
Equivariance of AugCLIP under Description Randomness
Using GPT-4V for description extraction introduces slight randomness across different seeds. To demonstrate near-equivalence quantitatively, we designed a randomness test on the attribute extraction process using the TEdBench dataset, varying five different seeds (examples in Appendix A. 2). ROUGE-L score among attributes extracted by different random seeds is 0.7590 (1 indicates maximum similarity), and the variance of AugCLIP in terms of human alignment is negligible, with a value of 0.0197. We also demonstrate in the Appendix that extracted attributes with different seeds do not vary.
@@
Our method, AugCLIP, adds additional computation time compared to $\mathrm{CLIP}_{\text{dir}}$. It requires extracting attributes via MLLM, and then fitting a hyperplane to distinguish source and target attributes. This process takes around 12.3 seconds for attribute generation and 0.15 seconds for score computation. This includes the non-reducible CLIP loading time (3.61 sec), which is required by other CLIP-based metrics, and the maximum attribute extraction time (8.62 sec). We will provide extracted attributes for all benchmark datasets to reduce this overhead, facilitating the evaluation process for future usage.
@@ Conclusion
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grants (RS-2023-00209060 A Study on Optimization and Network Interpretation Method for Large-Scale Machine Learning), Institute for Information & communications Technology Planning & Evaluation(IITP) grants (No.2022-0-00713 Meta-learning applicable to real-world problems, RS-2019-II190075 Artificial Intelligence Graduate School Program(KAIST)) funded by the Korea government(MSIT).
@@ A.3 Optimization Objective
Additionally, we visualize that CLIP features of source and target attributes can be separated by a linear hyperplane in 2D projected space using UMAP [21] in Fig. 8, in which randomly chosen subset of TEdBench samples are plotted. ‘S’ represents source attributes encoded into CLIP, while ‘T’ represents the target attributes. The line signifies the linear hyperplane $g(x)=\mathbf{W}^{T}x+b=0$ that separates the two classes, which are source and target. In AugCLIP, the linear hyperplane $g(x)=0$ is $d$-dimensional following the original dimension of CLIP, but to visualize in Fig. 8, the dimension shrinks into $d=2$ by UMAP fitting.